I. Anytown, suppose that one person is willing to pay $1,000 for relief In the year 2005, in from hay fever, another two are willing to pay S350; about five more are willing to pay $50; one is willing to pay $40; one is willing to pay $35; the rest of the town (another 75 people) are willing to pay $1. a. Draw the demand curve for hay fever relief in Anytown. (1 point) What is the potential total benefit from relief of hay fever if it is provided to everyone who asks? To everyone willing to pay S35 or more? (2 points) b. If the price of hay fever medication is S35, what is the quantity demanded? (1 point) c.
Answer #1

1200 Demand 1000 800 2 600 _Demand 400 200 10 12 Quantityb)

Total benefit = area under the demand curve and above the price level

i) when P = 0:

Total benefit = 1000 x 1 + 350 x 2 + 50 x 5 + 40 x 1 + 35 x 1 + 75 = 2100

ii) P = $ 35

Total benefit =  1000 x 1 + 350 x 2 + 50 x 5 + 40 x 1 + 35 x 1 = 2025

c)

Quantity demanded = 1 + 2 + 5 + 1 + 1 = 10